A 50g marble moving at 2.3m/s strikes a 21g marble at rest. Assume the collision is perfectly elastic and the marbles collide head-on. a) What is the speed of the first marble immediately after the collision? b) What is the speed of the second marble immediately after the collision?
Answer #1
v1 = (m1-m2)u1 + 2m2u2 /(m1+m2)
= 29*2.3 + 2*21*0 /(71)
= 0.94 m/s

(b)Byconservation of momentum, we can find that
v2 = 3.24 m/s
Answer #2
v1 = (m1-m2)u1 + 2m2u2 /(m1+m2)
= 29*2.3 + 2*21*0 /(71)
= 0.94 m/s

(b)Byconservation of momentum, we can find that
v2 = 3.24 m/s
Answer #3
v1 = (m1-m2)u1 + 2m2u2 /(m1+m2)
= 29*2.3 + 2*21*0 /(71)
= 0.94 m/s

(b)Byconservation of momentum, we can find that
v2 = 3.24 m/s